OpenStudy (anonymous):

To solve the equation 2sin(2x) = cosx, you should rewrite it as.......? A. cosx(4sinx - 1) = 0 B. 2sinx + cosx = 0 C. cosx(2sinx - 1) = 0

OpenStudy (solomonzelman):

I am sure you have seen this before, \(\large\color{black}{ \displaystyle \sin(\color{red}{\rm a}+\color{blue}{\rm b})=\sin(\color{red}{\rm a})\cos(\color{blue}{\rm b})+\sin(\color{blue}{\rm b})\cos(\color{red}{\rm a}) }\) what if your 'a' and 'b' are same? (such that x+x) ?

OpenStudy (amorfide):

sin2x = 2SinxCosx re write it to get 2(2sinxcosx)=cosx 4sinxcosx-cosx=0 cosx(4sinx - 1)=0

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sin(\color{red}{\rm x }+\color{blue}{\rm x})=\sin(\color{red}{\rm x})\cos(\color{blue}{\rm x})+\sin(\color{blue}{\rm x})\cos(\color{red}{\rm x})~~~~=\color{green}{2} \sin(\color{blue}{\rm x})\cos(\color{red}{\rm x})}\)

OpenStudy (solomonzelman):

is this making sense?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2\sin(2\color{red}{\rm x }) = \cos(\color{red}{\rm x })}\) re-writing \(\large\color{black}{ \displaystyle \color{green}{2} \sin(\color{blue}{\rm x})\cos(\color{red}{\rm x})=\cos(\color{red}{\rm x })}\) then subtract \(\large\color{black}{ \displaystyle \cos(\color{red}{\rm x })}\) from both sides, and factor on the left hand side to get the answer.

OpenStudy (anonymous):

Awesome, thanks Solomon! My classes are all online so sometimes it's harder to remember whether we've gone over something or not.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

yes?